# Gaussian consensus processes and their Lyapunov exponents

### v1

@@ 2.4 Proof of Theorem 2
The ratio whose logarithm appears in the above formula is simply the squared length of the projection of row $k$ of $M$ onto the orthogonal complement of the span of the first $k-1$ rows of $M$. Row $k$ is the sum of two independent vectors, namely $\beta e_{k}$ and $\frac{\sqrt{\alpha}(1-\beta)}{\sqrt{N}}G_{k}$, where $G_{k}$ is a mean zero Gaussian vector whose covariance matrix is $I_{N-1}$. Let $S_{k-1}$ be the span of the first $k-1$ rows of $M$, and let $\pi_{k-1}$ denote the orthogonal projection onto the orthogonal complement of $S_{k-1}$. Then we can decompose $G_{k}$ as the sum of three mutually orthogonal components $G_{k}=v_{1}+v_{2}+v_{3}$, where $v_{1}\in S_{k-1}$ and $v_{2}$ is in the span of $\pi_{k-1}(e_{k})$. Row $k$ is independent of $S_{k-1}$, so conditional on $S_{k-1}$, $v_{1}$, $v_{2}$ and $v_{3}$ are conditionally independent Gaussian vectors, and conditional on $S_{k-1}$, $\|v_{2}\|^{2}$ has a $\chi^{2}(1)$ distribution, and $\|v_{3}\|^{2}$ has a $\chi^{2}(N-k-1)$ distribution. Since these distributions do not depend on $S_{k-1}$, it is also holds unconditionally that $\|v_{2}\|^{2}$ and $\|v_{3}\|^{2}$ are independent with these two $\chi^{2}$ distributions. Finally, $v_{2}$ is parallel or antiparallel to $\pi_{k-1}(e_{k})$ with equal probability conditional on $\|v_{2}\|$. Recall that $\rho=\alpha(1-\beta)^{2}$. We obtain
@@
=$\stackrel{{\scriptstyle d}}{{=}}$ $\left(\frac{\sqrt{\rho}}{\sqrt{N}}X_{1}+\beta W_{k}\right)^{2}+\frac{\sqrt{\rho}}{\sqrt{N}}\sum_{i=1}^{N-k}X_{i}^{2}\,,$
@@
For $21\leq k\leq N-$ 2, note that $W_{k}$ strictly stochastically dominates $W_{k+1}$. To see this, observe that $W_{k}$ is also the distribution of $P_{k-1}(e_{N-1})$, since the law of $S_{k-1}$ is invariant under the orthogonal transformation which exchanges coordinates $k$ and $N-1$. Note that $P_{k}=P_{k}P_{k-1}$ so $P_{k}(e_{N-1})=P_{k}(P_{k-1}(e_{N-1})$. Define
@@
  • the sign of $X_{1}$ is independent of the absolute values $|X_{1}|,\dots,|X_{N}|$ and $\mathbb{P}(X_{1}<0)=1/2$,
@@
  • the events that $|X_{N-k}|=\max(|X_{1}|,\dots,|X_{N-k}|)$ and $\max(|X_{1}|,\dots,|X_{N-k}|)\geq 1$ are independent,
@@
$\mathbb{P}\left(E_{k}\right)$ $=\mathbb{P}(X_{1}<0,\ |X_{N-k}|\geq 1\;\text{ and }\;|X_{N-k}|=\max(|X_{1}|,\dots,|X_{N-k}|))>\frac{1}{2}\times 0.3\times\frac{1}{N-k}=\frac{0.15}{N-k}\,.$
@@
$\frac{V_{k+1}}{V_{k}}=1-\frac{\frac{\alpha(1-\beta)^{2}}{N}X_{N-k}^{2}}{\left(\beta W_{k}+\frac{\sqrt{\alpha}(1-\beta)}{\sqrt{N}}X_{1}\right)^{2}+\sum_{i=2}^{N-k}\frac{\alpha(1-\beta)^{2}}{N}X_{i}^{2}}\leq 1-\frac{\frac{\alpha(1-\beta)^{2}}{N}X_{N-k}^{2}}{\beta^{2}+\sum_{i=1}^{N-k}\frac{\alpha(1-\beta)^{2}}{N}X_{N-k}^{2}}\leq 1-\frac{1}{\frac{N\beta^{2}}{\alpha(1-\beta)^{2}}+N-k}.$
@@
for every $k$ in the range $21\leq k\leq N-$ 2. Notice that this lower bound on the gap between successive Lyapunov exponents is asymptotic to $0.15\alpha(1-\beta)^{2}/(N(N-k))$ as $\beta$ approaches $1$ from below. ∎
@@
Since $\mathrm{Cov}^{X}(t)$ is a non-negative definite symmetric matrix, its operator norm is its largest eigenvalue. Then the fact that $D_{1,1}>D_{2,2}$ implies that the right-hand side of equation 26 is dominated asymptotically by the contribution of the top-left entry of $D(t)$, and hence
@@
The limit above is a random $d$ by $d$ matrix of rank one (since $B_{1}$ has rank one), as long as $C\neq 0$. It remains to check that $C\neq 0$ almost surely. Indeed, $C$ could only be zero if $B_{1}V\overline{X}(0)=0$, but we assumed that the initial opinions are almost surely not all equal, which implies that $V\overline{X}(0)\neq 0$; then since $B_{1}$ is uniformly distributed on the unit sphere $S^{N-2}$, up to it sign ambiguity, the probability that $B_{1}$ lies in the left kernel of $V\overline{X}(0)$ is zero.
@@ 2.5 Some properties of the Lyapunov exponents $\lambda_{i}(\alpha,\beta,N)$
Proposition 3.
Suppose that $N$ is odd and $\beta>0$. Let $m=(N-1)/2$. Then
$\lambda_{1}(\alpha,\beta,N)$ $=\log\beta+\frac{1}{2}\left(\mathrm{Ei}_{1}\left(z\right)+\sum_{i=1}^{m-1}\frac{(-1)^{i}\,(i-1)!}{z^{i}}\left[e^{-z}-\binom{m-1}{i}\right]\right)$ (30)
@@
First, we will prove an auxiliary fact. Recall the definition of the function $\phi_{m}$ in equation (2).
Lemma 8.
For all $m>0$ we have
Proof.
where $\xi\sim\mathrm{Poi}(x)$. Note also that
The expression for $\phi_{m}^{\prime}(x)$ can be simplified further:
Iterating this formula $\ell$ times, we obtain that
Recalling that $\phi_{m}(0)=\psi(m)$, we obtain
To get the second expression for $\phi_{m}(x)$, note that
and thus
$\int_{0}^{1}\frac{(1-e^{-xs})(1-s)^{m-1}}{s}\mathrm{d}s$ $=-\int_{0}^{1}\left(\sum_{k=1}^{\infty}\frac{(-x)^{k}s^{k-1}}{k!}\right)(1-s)^{m-1}\mathrm{d}s=-\sum_{k=1}^{\infty}\frac{(-x)^{k}}{k\,\Gamma(k)}\int_{0}^{1}s^{k-1}(1-s)^{m-1}\mathrm{d}s$
$=-\sum_{k=1}^{\infty}\frac{(-x)^{k}}{k\,\Gamma(k)}\cdot\frac{\Gamma(k)\,\Gamma(m)}{\Gamma(m+k)}=-\sum_{k=1}^{\infty}\frac{(-x)^{k}}{k\cdot m(m+1)\dots(m+k-1)}.$
Proof of Proposition 3.
In case of an integer $m=(N-1)/2$, the expression obtained in (32) can be further simplified to
Let
Then
and
$=\sum_{j=1}^{m-1}\frac{(j-1)!}{(-x)^{j}}\sum_{i=j}^{m-1}\binom{i-1}{j-1}+\sum_{i=1}^{m-1}\frac{1}{i}+O(x)=\sum_{j=1}^{m-1}\frac{(j-1)!}{(-x)^{j}}\binom{m-1}{j}+\sum_{i=1}^{m-1}\frac{1}{i}+O(x)=B+\sum_{i=1}^{m-1}\frac{1}{i}+O(x)$
$\lim_{x\downarrow 0}(A-B)=\sum_{i=1}^{m-1}\frac{1}{i},$
which together with Theorem 1 yields the statement of the claim. ∎
Remark 6.
@@
Set $z=Ny/2$, where $y=\frac{\beta^{2}}{\rho}=\frac{\beta^{2}}{\alpha(1-\beta)^{2}}=O(1)$ and $y>0$. Recalling that $m=(N-1)/2$, from Theorem 1 we have
@@
where $\xi_{1},\dots,\xi_{N-1}$ are i.i.d. standard normal, independent of $X_{1},\dots,X_{N-2}$. Hence
@@
$\begin{array}[]{clrl}A_{0}&=\left\{|\xi_{3}^{2}+\dots+\xi_{N-1}^{2}-N|<N^{5/8}\right\},&A_{1}&=\left\{|\xi_{1}|<N^{1/8},\ |\xi_{2}|<N^{1/8}\right\},\\ B_{0}&=\left\{|X_{2}^{2}+\dots+\xi_{N-2}^{2}-N|<N^{5/8}\right\},&B_{1}&=\left\{|X_{1}|<N^{1/8}\right\},\end{array}$
@@
Hence, noting that the events in $A_{0}$ and $B_{0}$ deal with chi-squared random variables with $N-3$ degrees of freedom, and using the trivial bounds for standard normal random variables, we conclude (set $x=N^{1/4}$) that
@@
[…]
$+\mathbb{E}\left[\frac{\rho}{N}\left(\sum_{i=1}^{N-2}X_{i}^{2}-N\right)\mathbf{1}_{\mathscr{A}}\right]=-\frac{\rho\beta^{2}}{N}\cdot\frac{1+o(1)}{\rho+\beta^{2}}+0-\frac{2\rho+o(1)}{N}=-\frac{\rho}{N}\cdot\frac{2\rho+3\beta^{2}}{\rho+\beta^{2}}+o\left(\frac{1}{N}\right)$
@@
where $Y=\sum_{i=2}^{N-2}X_{i}^{2}-(N-3)$ is centred chi-squared random variable with $N-3$ degrees of freedom, using again the fact that $X_{1}$ is symmetric, and $\mathbb{E}\left[Y^{2}\mathbf{1}_{\mathscr{A}^{c}}\right]$ is exponentially small. Summarizing,
@@
[…]
@@
[…]
@@
Figure 5: $\alpha_{cr}$ as function of $N$ for $\beta=0$
@@
where $\rho=\alpha(1-\beta^{2})$. Let $\rho_{\mathrm{cr}}(\beta)=(1-\beta)^{2}\alpha_{\mathrm{cr}}(\beta)$ be the $\rho$ for which the displayed equality above holds for a given $\beta$. Since $\phi_{m}(x)>\phi_{m}(0)\equiv\psi(m)$ by Lemma 8, we get for $\beta>0$
@@
Lemma 12.
If $m\geq 1$, then as $x\to\infty$
$\phi_{m}(x)=\log x+\frac{m-1}{x}+o\left(\frac{1}{x}\right).$
Proof.
From the integral representation
$\phi_{m}(x)-\psi(m)$ $=\int_{0}^{1}\frac{(1-e^{-xs})(1-s)^{m-1}}{s}\mathrm{d}s$
$=\int_{0}^{1}\frac{1-e^{-xs}}{s}\mathrm{d}s-(m-1)\int_{0}^{1}1-e^{-xs}\mathrm{d}s+\int_{0}^{1}\frac{1-e^{-xs}}{s}\left[(1-s)^{m-1}-1+(m-1)s\right]\mathrm{d}s$
$=\gamma+\log x+\mathrm{Ei_{1}}(x)-(m-1)+\frac{(m-1)(1-e^{-x})}{x}+I$
where
$=\int_{0}^{1}\sum_{k=2}^{\infty}(-1)^{k}\binom{m-1}{k}s^{k-1}\mathrm{d}s-\int_{0}^{1}\frac{e^{-xs}}{s}\left[(1-s)^{m-1}-1+(m-1)s\right]\mathrm{d}s$
$=\sum_{k=2}^{\infty}\frac{(-1)^{k}}{k}\binom{m-1}{k}+R=(m-1)-\psi(m)-\gamma+R$
(from the Newton series for the digamma function) where
Note that $\sup_{s\in[0,1]}|g(s)|\leq m-1$ and by Peano’s formula
$g(s)=(m-1)(m-2)(1-\theta(s))^{m-3}\frac{s^{2}}{2!}.$
for some $0\leq\theta(s)\leq s$. Hence for $20\leq s\leq 1/$ 2 we have $|g(s)|\leq C_{m}s^{2}$ for some $C_{m}>0$. Consequently,
yielding
$\phi_{m}(x)-\log x$ $=\psi(m)+\gamma+\mathrm{Ei}_{1}(x)-(m-1)+\frac{m-1}{x}+I=\mathrm{Ei}_{1}(x)+\frac{m-1}{x}+O\left(\frac{1}{x^{2}}\right)=\frac{m-1}{x}+o(1/x)$
@@
$\alpha_{cr}(\beta)\sim\frac{2N}{N-3}\,\cdot\,\frac{1}{1-\beta}\quad\text{as $\beta\nearrow 1$.}$
@@
Make $z$ depend on $\beta$ by substituting $\alpha=\alpha_{cr}(\beta)$ into $z=\frac{N\beta^{2}}{2\alpha(1-\beta)^{2}}$. From the equality
$\lambda_{1}(\alpha_{cr}(\beta),\beta,N)=\log\beta+\frac{1}{2}\left[\phi_{\frac{N-1}{2}}(z)-\log z\right]=0$
@@
$0=2\log\beta+\frac{m-1}{z}(1+o(1)))=2\log\beta+\frac{(N-3)/2}{\frac{N\beta^{2}}{2\alpha_{cr}(\beta)(1-\beta)^{2}}}(1+o(1))$

### v2

@@ 2.4 Proof of Theorem 2
The ratio whose logarithm appears in the above formula is simply the squared length of the projection of row $k$ of $M$ onto the orthogonal complement of the span of the first k1k-1 rows of $M$. Row $k$ is the sum of two independent vectors, namely $\beta e_{k}$ and $\frac{\sqrt{\alpha}(1-\beta)}{\sqrt{N}}G_{k}$, where $G_{k}$ is a mean zero Gaussian vector whose covariance matrix is $I_{N-1}$. Let $S_{k-1}$ be the span of the first k1k-1 rows of $M$, and let $\pi_{k-1}$ denote the orthogonal projection onto the orthogonal complement of $S_{k-1}$. Consider the three mutually orthogonal subspaces $S_{k-1}$, $\langle\pi_{k-1}(e_{k})\rangle$, and $(S_{k-1}+\langle e_{k}\rangle)^{\perp}$, of dimensions k1k-1, 11 and Nk1N-k-1. By projecting $G_{k}$ onto these three subspaces, we may express $G_{k}$ as the sum of three mutually orthogonal vectors $G_{k}=v_{1}+v_{2}+v_{3}$, where $v_{1}\in S_{k-1}$ and $v_{2}$ is a scalar multiple of $\pi_{k-1}(e_{k})$. Row $k$ of $M$ is a Gaussian vector independent of $S_{k-1}$, with covariance a multiple of the identity. So conditional on $S_{k-1}$, the orthogonal summands $v_{1}$, $v_{2}$ and $v_{3}$ are conditionally independent Gaussian vectors. Moreover, conditional on $S_{k-1}$, $\|v_{2}\|^{2}$ has a $\chi^{2}(1)$ distribution, and $\|v_{3}\|^{2}$ has a $\chi^{2}(N-k-1)$ distribution, and these squared lengths are conditionally independent. Since these distributions do not depend on $S_{k-1}$, it also holds unconditionally that $\|v_{2}\|^{2}$ and $\|v_{3}\|^{2}$ are independent with these two $\chi^{2}$ distributions (even though $v_{2}$ and $v_{3}$ are dependent). Finally, $v_{2}$ is parallel or antiparallel to $\pi_{k-1}(e_{k})$ with equal probability conditional on $\|v_{2}\|$. Recall that $\rho=\alpha(1-\beta)^{2}$. We obtain
@@
$\stackrel{{\scriptstyle d}}{{=}}$ $\left(\frac{\sqrt{\rho}}{\sqrt{N}}X_{1}+\beta W_{k}\right)^{2}+{\frac{\rho}{N}}\sum_{i=1}^{N-k}X_{i}^{2}\,$,
@@
For $1\leq k\leq N-2$, note that $W_{k}$ strictly stochastically dominates $W_{k+1}$. To see this, observe that $W_{k}$ is also the distribution of $P_{k-1}(e_{N-1})$, since the law of $S_{k-1}$ is invariant under the orthogonal transformation which exchanges coordinates $k$ and N1N-1. Note that $P_{k}=P_{k}P_{k-1}$ so $P_{k}(e_{N-1})=P_{k}(P_{k-1}(e_{N-1}))$. Define
@@
  • $\mathbb{P}(X_{1}<0)=1/2$ and the sign of $X_{1}$ is independent of the absolute values $|X_{1}|,\dots,|X_{N}|$,
@@
  • the events $|X_{N-k}|=\max(|X_{1}|,\dots,|X_{N-k}|)$ and $\max(|X_{1}|,\dots,|X_{N-k}|)\geq 1$ are independent,
@@
$\mathbb{P}\left(E_{k}\right)$ $=\mathbb{P}(X_{1}<0,\ |X_{N-k}|\geq 1\;\text{ and }\;|X_{N-k}|=\max(|X_{1}|,\dots,|X_{N-k}|))$
$>\frac{1}{2}\times 0.3\times\frac{1}{N-k}=\frac{0.15}{N-k}\,.$
@@
$\frac{V_{k+1}}{V_{k}}$ $=1-\frac{\frac{\alpha(1-\beta)^{2}}{N}X_{N-k}^{2}}{\left(\beta W_{k}+\frac{\sqrt{\alpha}(1-\beta)}{\sqrt{N}}X_{1}\right)^{2}+\sum_{i=2}^{N-k}\frac{\alpha(1-\beta)^{2}}{N}X_{i}^{2}}\leq 1-\frac{\frac{\alpha(1-\beta)^{2}}{N}X_{N-k}^{2}}{\beta^{2}+\sum_{i=1}^{N-k}\frac{\alpha(1-\beta)^{2}}{N}X_{N-k}^{2}}$
$\leq 1-\frac{1}{\frac{N\beta^{2}}{\alpha(1-\beta)^{2}}+N-k}.$
@@
for every $k$ in the range $1\leq k\leq N-2$. Notice that this lower bound on the gap between successive Lyapunov exponents is asymptotic to $0.15\alpha(1-\beta)^{2}/(N(N-k))$ as $\beta$ approaches $1$ from below. ∎
@@
Since $\mathrm{Cov}^{X}(t)$ is a non-negative definite symmetric matrix, its operator norm is its largest eigenvalue. Then the fact that $D_{1,1}>D_{2,2}$ implies that the right-hand side of equation (27) is dominated asymptotically by the contribution of the top-left entry of $D(t)$, and hence
@@
The limit above is a random $d$ by $d$ matrix of rank one (since $B_{1}$ has rank one), as long as $C\neq 0$. It remains to check that $C\neq 0$ almost surely. Indeed, $C$ could only be zero if $B_{1}V\overline{X}(0)=0$, but we assumed that the initial opinions are almost surely not all equal, which implies that $V\overline{X}(0)\neq 0$; then since $B_{1}$ is uniformly distributed on the unit sphere $\mathbb{S}^{N-2}$, up to its sign ambiguity, the probability that $B_{1}$ lies in the left kernel of $V\overline{X}(0)$ is zero.
@@ 2.5 Some properties of the Lyapunov exponents $\lambda_{i}(\alpha,\beta,N)$
$\mathrm{Ei}_{1}(x)=\int_{1}^{\infty}\frac{e^{-xu}}{u}du=\int_{0}^{x}\frac{1-e^{-u}}{u}du-\gamma-\log(x).$
Proposition 3.
Suppose that $N$ is odd and $\beta>0$. Let m=(N1)/2m=(N-1)/2. Then
@@
Set $z=Ny/2$, where $y=\frac{\beta^{2}}{\rho}=\frac{\beta^{2}}{\alpha(1-\beta)^{2}}=O(1)$ and $y>0$. Recalling that m=(N1)/2m=(N-1)/2, from Theorem 1 we have
@@
where $\xi_{1},\dots,\xi_{N-1}$ are i.i.d. standard normal, independent of $X_{1},\dots,X_{N-2}$. Indeed, the first row of $M(1)$ can be written as $\left(\beta+\sqrt{\frac{\rho}{N}}\xi_{1},\sqrt{\frac{\rho}{N}}\xi_{2},\dots,\sqrt{\frac{\rho}{N}}\xi_{N-1}\right)$, and the projection of $e_{2}$ on the linear space spanned by this vector can be obtained by minimizing
$\left[k\left(\beta+\sqrt{\frac{\rho}{N}}\xi_{1}\right)\right]^{2}+\left[k\sqrt{\frac{\rho}{N}}\xi_{2}-1\right]^{2}+\left[k\sqrt{\frac{\rho}{N}}\xi_{3}\right]^{2}+\dots+\left[k\sqrt{\frac{\rho}{N}}\xi_{N-1}\right]^{2}$
with respect to $k\in\mathbb{R}$. Then, using the Pythagoras theorem, we obtain (35). Hence
@@
$\begin{array}[]{clrl}A_{0}&=\left\{|\xi_{3}^{2}+\dots+\xi_{N-1}^{2}-N|<N^{5/8}\right\},&A_{1}&=\left\{|\xi_{1}|<N^{1/8},\ |\xi_{2}|<N^{1/8}\right\},\\ B_{0}&=\left\{|X_{2}^{2}+\dots+X_{N-2}^{2}-N|<N^{5/8}\right\},&B_{1}&=\left\{|X_{1}|<N^{1/8}\right\},\end{array}$
@@
Hence, noting that the events in $A_{0}$ and $B_{0}$ deal with chi-squared random variables with N3N-3 degrees of freedom, and using the trivial bounds for standard normal random variables, we conclude (set $x=N^{1/4}$) that
@@
$(\rho+\beta^{2})\mathbb{E}[\eta\mathbf{1}_{\mathscr{A}}]$ $=-\frac{\rho\beta^{2}}{N}\mathbb{E}\left[\frac{\xi_{2}^{2}\,\mathbf{1}_{\mathscr{A}}}{(\beta+\sqrt{\rho/N}\xi_{1})^{2}+\rho\xi_{2}^{2}/N+\rho(\xi_{3}^{2}+\dots+\xi_{N-1}^{2})/N}\right]$
$+\mathbb{E}\left[2\beta W_{2}\frac{\sqrt{\rho}}{\sqrt{N}}X_{1}\mathbf{1}_{\mathscr{A}}\right]+\mathbb{E}\left[\frac{\rho}{N}\left(\sum_{i=1}^{N-2}X_{i}^{2}-N\right)\mathbf{1}_{\mathscr{A}}\right]$
$=-\frac{\rho\beta^{2}}{N}\cdot\frac{1+o(1)}{\rho+\beta^{2}}+0-\frac{2\rho+o(1)}{N}=-\frac{\rho}{N}\cdot\frac{2\rho+3\beta^{2}}{\rho+\beta^{2}}+o\left(\frac{1}{N}\right)$
@@
where $Y=\sum_{i=2}^{N-2}X_{i}^{2}-(N-3)$ is centred chi-squared random variable with N3N-3 degrees of freedom, using again the fact that $X_{1}$ is symmetric, and $\mathbb{E}\left[Y^{2}\mathbf{1}_{\mathscr{A}^{c}}\right]$ is exponentially small. Summarizing,
@@
$\left|\mathbb{E}\left[\log(Q)\,\mathbf{1}_{\mathscr{A}^{c}}\right]\right|$ $\leq\sqrt{\mathbb{E}\log^{2}(Q)}\sqrt{\mathbb{E}\left[\mathbf{1}_{\mathscr{A}^{c}}\right]}=\sqrt{\mathbb{E}\log^{2}(Q)}\cdot\sqrt{\mathbb{P}(\mathscr{A}^{c})}$
@@
$\mathbb{E}\left[\log^{2}(Q)\,\mathbf{1}_{Q\geq 1}\right]$ $\leq\mathbb{E}\left[\log^{2}(Q)\,\mathbf{1}_{Q\geq 1}\right]\leq\mathbb{E}[Q^{2}]=O(1);$
$\mathbb{E}\left[\log^{2}(Q)\,\mathbf{1}_{Q<1}\right]$ $\leq\mathbb{E}\left[Q^{-1}\,\mathbf{1}_{Q<1}\right]\leq\mathbb{E}[Q^{-1}]\leq\frac{N}{\rho}\mathbb{E}\left[\frac{1}{X_{2}^{2}+\dots+X_{N-2}^{2}}\right]=\frac{N}{\rho(N-5)}=O(1),$
@@
Figure 3: $\alpha_{\mathrm{cr}}$ as function of $N$ for $\beta=0$
@@
where $\rho=\alpha(1-\beta)^{2}$. Let $\rho_{\mathrm{cr}}(\beta)=(1-\beta)^{2}\alpha_{\mathrm{cr}}(\beta)$ be the $\rho$ for which the displayed equality above holds for a given $\beta$. Since $\phi_{m}(x)>\phi_{m}(0)\equiv\psi(m)$ by Lemma 13, we get for $\beta>0$
@@
$\alpha_{\mathrm{cr}}(\beta)\sim\frac{2N}{N-3}\,\cdot\,\frac{1}{1-\beta}\quad\text{as $\beta\nearrow 1$.}$
@@
Make $z$ depend on $\beta$ by substituting $\alpha=\alpha_{\mathrm{cr}}(\beta)$ into $z=\frac{N\beta^{2}}{2\alpha(1-\beta)^{2}}$. From the equality
$\lambda_{1}(\alpha_{\mathrm{cr}}(\beta),\beta,N)=\log\beta+\frac{1}{2}\left[\phi_{\frac{N-1}{2}}(z)-\log z\right]=0$
@@
$0=2\log\beta+\frac{m-1}{z}(1+o(1)))=2\log\beta+\frac{(N-3)/2}{\frac{N\beta^{2}}{2\alpha_{\mathrm{cr}}(\beta)(1-\beta)^{2}}}(1+o(1))$